Suppose you like to keep a jar of change on your desk. Currently, the jar contains the following:

10 Pennies

7 Dimes

18 Nickels

27 Quarters

What is the probability that you reach into the jar and randomly grab a penny and then, without replacement, a dime? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Probability that we get a penny in first attempt is given as no of pennies divided by total number of changes.

Thus required probability is 10/(10+7+18+27)=10/62=0.1613

Next we don't replace the change and then probability of getting a dime becomes 7/61=0.1147 where numerator is total number of dimes and denominator is total number of changes remaining.

That's the answer.

Now happening of both events together can be given as (10/62)*(7/61)=.0185.